OpenStudy (anonymous):

Write the equation of the line that passes through (3, 6) and (0, –3) in slope-intercept form. y = 3x – 6 y = 3x + 3 y = 3x + 6 y = 3x – 3

OpenStudy (saifoo.khan):

D

OpenStudy (anonymous):

D. Slope = (y2-y1)/(x2-x1) (-3-6)/(0-3) = 3 = slope y = mx +b y = 3x + b 6 =3(3) + b 6 = 9 +b -9=-9 -3=b